# Optimal Conformal Prediction, E-values, Fuzzy Prediction Sets and Subsequent Decisions

### v1

@@ 4.6 Optimal classical conformal
In Theorem 8, we show how our optimality framework nests classical conformal prediction as a special case if we choose the Neyman–Pearson utility function $U_{\alpha}^{\textnormal{NP}}:x\mapsto x\wedge 1/\alpha$, $\alpha\in(0,1)$. This shows our result nests classical conformal prediction for any non-conformity score that is a monotone function of $\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}=(n+1)\times Q^{Z_{n+1}|O(z^{n+1})}$.
@@
$\varepsilon^{\textnormal{NP}}(z^{n+1})=\begin{cases}1/\alpha,&\textnormal{ if }\textnormal{LR}^{Z_{n+1}|O^{\prime}}(z_{n+1})>c_{\alpha}^{O^{\prime}},\\ k^{O^{\prime}},\hskip-11.38092pt&\textnormal{ if }\textnormal{LR}^{Z_{n+1}|O^{\prime}}(z_{n+1})=c_{\alpha}^{O^{\prime}},\\ 0,&\textnormal{ if }\textnormal{LR}^{Z_{n+1}|O^{\prime}}(z_{n+1})<c_{\alpha}^{O^{\prime}},\end{cases}$
where $c_{\alpha}^{O^{\prime}}$ is the $\alpha$ upper-quantile of $\textnormal{LR}^{|O^{\prime}}(z^{n+1})$ under $\textnormal{Unif}[Z^{n+1}]$, and $k^{O^{\prime}}$ is some orbit-dependent constant that ensures $\varepsilon^{\textnormal{NP}}(Z^{n+1})$ is an exact e-value. This e-value is optimal uniformly in any mixture over orbits, and in any monotone transformation of the likelihood ratio.
To make this more interpretable, we provide Corollary 1, which combines Theorem 8 with our result for the i.i.d. setting in Proposition 2 for minimizing the expected confidence set
$\mathrm{E}^{P_{*}^{Z^{n}}}[\mu(C_{\alpha}^{Z^{n}})].$
Using $f=dP_{*}^{Z_{n+1}}/d\mu$ to denote the density of the true distribution, Corollary 1 shows that classical conformal prediction is optimal in this sense when the conformity score $s$ is chosen as a monotone function of the density $f$. Conversely, given a conformity score $s$ and measure $\mu$, classical conformal prediction is optimal uniformly in the class of true distributions for which $s$ is a monotone function of the corresponding density $f$.
@@
Suppose $P_{*}^{Z^{n+1}}$ is i.i.d., and let $s:\mathcal{Z}\to\mathbb{R}_{+}$ denote a conformity score. Then, classical conformal prediction minimizes the expected confidence set size for any measure $\mu$ and true distribution $P_{*}^{Z^{n+1}}$ that satisfies
$s=m\left(\frac{dP_{*}^{Z_{n+1}}}{d\mu}\right),$
where $m:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is some monotone function.
5 Conformal certified decisions
In this section, we show how (fuzzy) prediction confidence sets may be used in subsequent decision making.
We consider a decision maker who is faced with selecting a decision $d$ in a decision space $\mathcal{D}$ while facing uncertainty captured by a random variable $Z_{n+1}$ in the outcome space $\mathcal{Z}$. When selecting the decision $d$, they would ideally like to minimize the ‘oracle loss’ $L:\mathcal{D}\times\mathcal{Z}\to\mathbb{R}$ associated with the decision under the true outcome of $Z_{n+1}$: $\min_{d\in\mathcal{D}}L(d,Z_{n+1})$. Unfortunately, the outcome of $Z_{n+1}$ is not yet available at the time of decision making. Instead, the decision maker relies on a (fuzzy) confidence set for $Z_{n+1}$ to aid them in their decision making process.
5.1 Confidence sets and as-if decisions
We begin by presenting the case where the decision maker receives a conventional non-fuzzy confidence set, before extending to fuzzy confidence sets. For these non-fuzzy confidence sets, this was recently developed byAndrews and Chen (2025) and Kiyani et al. (2025).
Suppose the decision maker is given a valid level $(1-\alpha)$ confidence set $C_{\alpha}^{Z^{n}}\subseteq\mathcal{Z}$ for $Z_{n+1}$ under $\mathcal{P}$. We define the as-if decision $\delta(C_{\alpha}^{Z^{n}})$, as the decision that minimizes the worst-case loss over all possible realizations of $Z_{n+1}$ in the confidence set $C_{\alpha}^{Z^{n}}$:
@@ 5.1 Confidence sets and as-if decisions
where we assume the minimum exists (Manski, 2021). Note that if the confidence set happens to only contain the true future realization of $Z_{n+1}$, then this action would minimize the oracle loss.
Kiyani et al. (2025) and Andrews and Chen (2025) show that this as-if decision $\delta(C_{\alpha}^{Z^{n}})$ enjoys a $(1-\alpha)$-confidence lower bound on the to-be-realized loss:
@@
5.2 Fuzzy confidence sets
We now move to fuzzy confidence sets, and showcase two distinct ways in which they generalize the as-if decision making framework: as-if decisions under the post-hoc selection of the confidence level $\alpha$, or, more generally, minimax decisions over a loss function weighted by the fuzzy confidence set.
Minimax decisions over an evidence-weighted loss function have previously been for classical confidence set on parameter spaces under the name ‘E-posterior’ and ‘E-certification’ (Grünwald, 2023b; Andrews and Chen, 2025). While post-hoc level selection has been explored in the context of conformal prediction before (Koning, 2025a; Gauthier et al., 2025), the idea of post-hoc as-if decisions is novel, to the best of our knowledge. Moreover, we are also the first to consider such weighted decisions outside of a parameter space.
@@ 5.2.1 Post-hoc level as-if decisions
A downside of subsequent decisions with non-fuzzy confidence sets, as presented in Section 5.1, is that the decision maker must pre-specify a single confidence level $\alpha$. In this section, we show that a fuzzy confidence set $\mathcal{E}^{Z^{n}}$ enables a decision maker to choose among as-if decisions at each confidence level $\alpha$, and select the decision with the desired confidence level $\alpha$ in an arbitrary data-dependent fashion. Overall, this still provides a post-hoc $\alpha$ confidence guarantee.
@@
$C_{\alpha}^{Z^{n}}$ $=\left\{z\in\mathcal{Z}:\mathcal{E}^{Z^{n}}(z)<1/\alpha\right\}.$
For each sublevel set, we then construct the as-if decision
$\delta(C_{\alpha}^{Z^{n}})=\operatorname*{arg\,min}_{d\in\mathcal{D}}\sup_{z\in C_{\alpha}^{Z^{n}}}L(d,z).$
This produces a collection $(\delta(C_{\alpha}^{Z^{n}}))_{\alpha>0}$ of possible decisions. A decision maker may then select the desired level of confidence $\widetilde{\alpha}$ in an arbitrary data-dependent fashion. The resulting associated decision then satisfies the the confidence guarantee
$\mathrm{E}_{\widetilde{\alpha}}^{P}\left[\frac{P\left(L(\delta(C_{\widetilde{\alpha}}^{Z^{n}},z)\geq R(\delta(C_{\widetilde{\alpha}}^{Z^{n}}))\mid\widetilde{\alpha}\right)}{\widetilde{\alpha}}\right]\leq 1,$ (19)
@@
As a traditional (non-fuzzy) level $\alpha$ confidence set may be seen as a special binary fuzzy confidence set, we may technically also use data-dependent levels with traditional confidence sets. However, we then find that if we select $\widetilde{\alpha}>\alpha$ yields the same confidence set as $\alpha$ but with a weaker guarantee $\widetilde{\alpha}$, and if $\widetilde{\alpha}<\alpha$ then the resulting post-hoc confidence set is the entire sample space. In this sense, non-fuzzy level $\alpha$ confidence sets effectively forces the choice $\widetilde{\alpha}=\alpha$.
@@ 5.2.2 Weighted loss
Instead of producing a collection of possible as-if decisions under a post-hoc level guarantee, it is also possible to weigh our loss by the values of the fuzzy confidence set.
@@
Such weighted loss bounds have been considered for confidence sets on parameter spaces by Grünwald (2023b) and Andrews and Chen (2025), but to the best of our knowledge we are the first to consider them on a sample space.
@@ 5.3 Connecting as-if and weighted
While the weighted loss-approach for fuzzy confidence sets in Section 5.2.2 and the as-if decision from in Section 5.1 may seem distinct, we show that non-fuzzy approach may be viewed as a special limiting case of the fuzzy angle. This relationship was not yet established before; Andrews and Chen (2025) describes both approaches as distinct paradigms. The connection is not obvious, as it relies on our observation that the E-posterior coined by Grünwald (2023b) is equivalent to a fuzzy confidence set, and so a generalization of a confidence set.
Let $\gamma\in(0,1)$, and define the following fuzzy confidence set for every $z\in\mathcal{Z}$:
$\mathcal{E}^{Z^{n}}(z):=\gamma\frac{1}{\alpha}\mathbb{I}\{z\notin\mathcal{C}_{\alpha}^{Z^{n}}\}+(1-\gamma)\frac{1}{\alpha}\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}.$ (20)
Note that when $\gamma\to 1$, then $\mathcal{E}$ converges to non-fuzzy confidence set. Moreover, this is indeed a fuzzy confidence set, since
$\mathrm{E}^{P}[\mathcal{E}^{Z^{n}}(Z_{n+1})]$
$=\gamma\frac{1}{\alpha}P(Z_{n+1}\notin\mathcal{C}_{\alpha}^{Z^{n}})$
$+(1-\gamma)\frac{1}{\alpha}(1-P(Z_{n+1}\not\in\mathcal{C}_{\alpha}^{Z^{n}}))$
$=\frac{1}{\alpha}P(Z_{n+1}\notin\mathcal{C}_{\alpha}^{Z^{n}})\leq 1.$
@@
$\sup_{z}\frac{L(d,z)}{\mathcal{E}^{Z^{n}}(z)}$
$=\sup_{z}\frac{L(d,z)}{\gamma\frac{1}{\alpha}\mathbb{I}\{z\notin\mathcal{C}_{\alpha}^{Z^{n}}\}+(1-\gamma)\frac{1}{\alpha}\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}}$
$=\sup_{z}\frac{\alpha}{\gamma}\mathbb{I}\{z\not\in\mathcal{C}_{\alpha}^{Z^{n}}\}L(d,z)+\frac{\alpha}{1-\gamma}\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}L(d,z).$
@@
$\sup_{z}\frac{\alpha}{1-\gamma}\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}L(d,z)=\frac{\alpha}{1-\gamma}\sup_{z\in\mathcal{C}_{\alpha}^{Z^{n}}}L(d,z),$
@@
which is the risk function we minimize for non-fuzzy confidence sets in (18), up to the constant $\alpha/(1-\gamma)$ that is irrelevant for the optimization. This shows that the non-fuzzy ‘as-if’ decision can indeed be viewed as a limiting case of the fuzzy procedure.

### v2

@@ 5.5 Optimal classical conformal
In Theorem 8, we show how our optimality framework nests classical conformal prediction as a special case if we choose the Neyman–Pearson utility function $U_{\alpha}^{\textnormal{NP}}:x\mapsto x\wedge 1/\alpha$, $\alpha\in(0,1)$. This nests classical conformal prediction for any non-conformity score that is a monotone function of $\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}=(n+1)\times Q^{Z_{n+1}|O(z^{n+1})}$.
@@
$\mathcal{E}^{\textnormal{NP}}(z^{n+1})=\begin{cases}1/\alpha,&\textnormal{ if }\textnormal{LR}^{Z_{n+1}|O^{\prime}}(z_{n+1})>c_{\alpha}^{O^{\prime}},\\ k^{O^{\prime}},\hskip-11.38092pt&\textnormal{ if }\textnormal{LR}^{Z_{n+1}|O^{\prime}}(z_{n+1})=c_{\alpha}^{O^{\prime}},\\ 0,&\textnormal{ if }\textnormal{LR}^{Z_{n+1}|O^{\prime}}(z_{n+1})<c_{\alpha}^{O^{\prime}},\end{cases}$
where $c_{\alpha}^{O^{\prime}}$ is the $\alpha$ upper-quantile of $\textnormal{LR}^{|O^{\prime}}(z^{n+1})$ under $\textnormal{Unif}[Z^{n+1}]$, and $k^{O^{\prime}}$ is some orbit-dependent constant that ensures $\mathcal{E}^{\textnormal{NP}}(Z^{n+1})$ is an exact e-value. This e-value is optimal uniformly in any mixture over orbits, and in any monotone transformation of the likelihood ratio.
To make this more interpretable, we provide Corollary 2, which combines Theorem 8 with our result for the i.i.d. setting in Proposition 2. Here, we consider an i.i.d. estimator $\widehat{P}^{Z^{n}}=(\widehat{P}^{Z_{1}})^{n}$ for $P_{*}^{Z^{n}}$, and minimize the expected measure of the prediction set
$\mathrm{E}^{\widehat{P}^{Z^{n}}}[\mu(C_{\alpha}^{Z^{n}})].$
Using $f=d\widehat{P}^{Z_{1}}/d\mu$ to denote the density, Corollary 2 shows that classical conformal prediction is optimal in this sense when the conformity score $s$ is chosen as a monotone function of the density $f$. Conversely, given a conformity score $s$ and measure $\mu$, classical conformal prediction is optimal uniformly in the class of true distributions for which $s$ is a monotone function of the corresponding density $f$.
@@
Suppose $\widehat{P}^{Z^{n}}=(\widehat{P}^{Z_{1}})^{n}$, and let $s:\mathcal{Z}\to\mathbb{R}_{+}$ denote a conformity score. Then, classical conformal prediction minimizes the expected prediction set size for any $\mu$ and $\widehat{P}^{Z_{1}}$ that satisfy
$s=m\left(\frac{d\widehat{P}^{Z_{1}}}{d\mu}\right),$
where $m:[0,\infty]\to[-\infty,\infty]$ is some strictly monotone function.
5.6 Clipping, capping and dampening
In this section, we cover some practical tools for designing utility functions.
$\mathcal{E}(z^{n+1})=\left(\lambda\textnormal{LR}^{Z_{n+1}|O(z^{n+1})}(z_{n+1})\right)\vee b,$ (14)
where $\lambda$ is some normalization constant that ensures $\mathcal{E}$ is an exact e-value. This may be interpreted as ‘clipping’ the likelihood ratio from below, and simultaneously shrinking it by $\lambda$ to ensure it remains a valid e-value. Alternatively, it can be viewed as choosing a utility function with $U(x)=$-$\infty$ on $[0,b)$.
Analogously, in case we are not interested in evidence beyond some value $c\geq 0$, we may also cap our e-values below some upper bound $c\geq 1$. The resulting optimizer is analogous to (14), with $\vee b$ is replaced by $\wedge c$ (koning2025continuoustestingunifyingtests). Capping may also be incorporated in the utility function, and appears in the Neyman–Pearson utility function $U(x)=x\wedge 1/\alpha$.
An alternative strategy to ‘dampen’ e-values is introduced by grunwald2023posterior. He proposes to take an e-value $\mathcal{E}$ and dampen it to the e-value $b+(1-b)\mathcal{E}$ that may be interpreted as a $b$-mixture of the e-value and the constant 1. If $\mathcal{E}$ is the log-utility maximizer, then we find $b+(1-b)\mathcal{E}$ implicitly maximizes the utility function
$U(\mathcal{E})=\log([\mathcal{E}-b]\vee 0).$
6 Certifying subsequent decisions
In this section, we show how (fuzzy) prediction sets may be used in subsequent decision making.
We consider a decision maker who is faced with selecting a decision $d$ in a decision space $\mathcal{D}$ while facing uncertainty captured by a random variable $Z_{n+1}$ in the outcome space $\mathcal{Z}$. When selecting the decision $d$, they would ideally like to minimize the ‘oracle loss’ $L:\mathcal{D}\times\mathcal{Z}\to\mathbb{R}$ associated with the decision under the true outcome of $Z_{n+1}$: $\min_{d\in\mathcal{D}}L(d,Z_{n+1})$. Unfortunately, the outcome of $Z_{n+1}$ is not yet available at the time of decision making. Instead, the decision maker relies on a (fuzzy) prediction set for $Z_{n+1}$ to aid them in their decision making process.
6.1 Prediction sets and as-if decisions
We begin by presenting the case where the decision maker receives a conventional non-fuzzy prediction set, before extending to fuzzy prediction sets. For these non-fuzzy prediction sets, this was recently developed by andrews2025certified and kiyani2025decision.
Suppose the decision maker is given a valid level $(1-\alpha)$ prediction set $C_{\alpha}^{Z^{n}}\subseteq\mathcal{Z}$ for $Z_{n+1}$ under $\mathcal{P}$. We define the as-if decision $\delta(C_{\alpha}^{Z^{n}})$, that pretends as if $Z_{n+1}\in C_{\alpha}$. That is, it minimizes the worst-case loss over all possible realizations of $Z_{n+1}$ in the prediction set $C_{\alpha}^{Z^{n}}$:
@@ 6.1 Prediction sets and as-if decisions
where we assume the minimum exists (manski2021econometrics). Note that if the prediction set happens to only contain the true future realization of $Z_{n+1}$, then this action would minimize the oracle loss.
kiyani2025decision and andrews2025certified show that this as-if decision $\delta(C_{\alpha}^{Z^{n}})$ enjoys a $(1-\alpha)$-confidence lower bound on the to-be-realized loss:
@@
6.2 Fuzzy prediction sets
We now move to fuzzy prediction sets, and showcase two distinct ways in which they generalize the as-if decision making framework: as-if decisions under the post-hoc selection of the confidence level $\alpha$ and minimax decisions over a loss function weighted by the fuzzy prediction set.
Minimax decisions over an evidence-weighted loss function have previously been for fuzzy confidence sets on parameter spaces under the name ‘E-posterior’ and ‘E-certification’ (grunwald2023posterior; andrews2025certified). While post-hoc level selection has been explored in the context of conformal prediction before (gauthier2025values), the idea of post-hoc as-if decisions is novel, to the best of our knowledge. Moreover, we are also the first to consider such weighted decisions outside of a parameter space.
@@ 6.2.1 Post-hoc level as-if decisions
A downside of subsequent decisions with non-fuzzy prediction sets, as presented in Section 6.1, is that the decision maker must pre-specify a single confidence level $\alpha$. In this section, we show that a fuzzy prediction set $\mathcal{E}^{Z^{n}}$ enables a decision maker to choose among as-if decisions at each confidence level $\alpha$, and select the decision with the desired confidence level $\alpha$ in an arbitrary data-dependent fashion. Overall, this still provides a post-hoc $\alpha$ confidence guarantee.
@@
$\overline{C}_{\alpha}^{Z^{n}}$ $=\left\{z\in\mathcal{Z}:\mathcal{E}^{Z^{n}}(z)<1/\alpha\right\}.$
For each sublevel set $\overline{C}_{\alpha}$, we then construct the as-if decision
$\delta(\overline{C}_{\alpha}^{Z^{n}})=\operatorname*{arg\,min}_{d\in\mathcal{D}}\sup_{z\in\overline{C}_{\alpha}^{Z^{n}}}L(d,z).$
This produces a collection $(\delta(\overline{C}_{\alpha}^{Z^{n}}))_{\alpha>0}$ of possible decisions. A decision maker may then select the desired level of confidence $\widetilde{\alpha}$ in an arbitrary data-dependent fashion. The resulting associated decision then satisfies the confidence guarantee
$\mathrm{E}_{\widetilde{\alpha}}^{P}\left[\frac{P\left[L(\delta(\overline{C}_{\widetilde{\alpha}}^{Z^{n}}),z)\geq R(\delta(\overline{C}_{\widetilde{\alpha}}^{Z^{n}}))\mid\widetilde{\alpha}\right]}{\widetilde{\alpha}}\right]\leq 1,$ (16)
@@
As a traditional (non-fuzzy) level $\alpha$ prediction set may be seen as a special binary fuzzy prediction set, we may technically also use data-dependent levels with traditional prediction sets. However, we then find that if we select $\widetilde{\alpha}>\alpha$ this yields the same prediction set as $\alpha$ but with a weaker guarantee $\widetilde{\alpha}$, and if $\widetilde{\alpha}<\alpha$ then the resulting prediction set is the entire sample space. In this sense, a non-fuzzy level $\alpha$ prediction set effectively forces the choice $\widetilde{\alpha}=\alpha$ as the only sensible option.
@@ 6.2.2 Weighted loss
Instead of producing a collection of possible as-if decisions under a post-hoc level guarantee, it is also possible to weight our loss by the values of the fuzzy prediction set.
@@ 6.3 Connecting as-if and weighted
While the weighted loss approach for fuzzy prediction sets in Section 6.2.2 and the as-if decision approach from Section 6.1 may seem distinct, we show that as-if approach may be viewed as a special limiting case of the weighted loss approach. This relationship was not yet established before; andrews2025certified describe both approaches as distinct paradigms. The connection is not obvious, as it relies on our observation that the E-posterior coined by grunwald2023posterior is equivalent to a fuzzy confidence set, and so a generalization of a confidence set.
Let $\gamma\in(0,1)$, and define the following fuzzy prediction set for every $z_{n+1}\in\mathcal{Z}$:
$\mathcal{E}^{Z^{n}}(z_{n+1}):=\gamma\frac{1}{\alpha}\mathbb{I}\{z_{n+1}\notin\mathcal{C}_{\alpha}^{Z^{n}}\}+(1-\gamma)\mathbb{I}\{z_{n+1}\in\mathcal{C}_{\alpha}^{Z^{n}}\}.$ (17)
Note that $\mathcal{E}$ converges to a non-fuzzy prediction set when $\gamma\to 1$. Moreover, this is indeed an e-value because it is dominated by the mixture $\gamma\frac{1}{\alpha}\mathbb{I}\{Z_{n+1}\notin\mathcal{C}_{\alpha}^{Z^{n}}\}+(1-\gamma)$ of a valid e-value $\frac{1}{\alpha}\mathbb{I}\{Z_{n+1}\notin\mathcal{C}_{\alpha}^{Z^{n}}\}$ and the constant 1, and a mixture of valid e-values is itself a valid e-value.
@@
$\sup_{z}\frac{L(d,z)}{\mathcal{E}^{Z^{n}}(z)}=\sup_{z}\frac{L(d,z)}{\gamma\frac{1}{\alpha}\mathbb{I}\{z\notin\mathcal{C}_{\alpha}^{Z^{n}}\}+(1-\gamma)\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}}$
$=\sup_{z}\left\{\frac{\alpha}{\gamma}\mathbb{I}\{z\not\in\mathcal{C}_{\alpha}^{Z^{n}}\}L(d,z)+\frac{1}{1-\gamma}\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}L(d,z)\right\}.$
@@
$\sup_{z}\frac{1}{1-\gamma}\mathbb{I}\{z\in\mathcal{C}_{\alpha}^{Z^{n}}\}L(d,z)=\frac{1}{1-\gamma}\sup_{z\in\mathcal{C}_{\alpha}^{Z^{n}}}L(d,z),$
@@
which is the risk function we minimize for non-fuzzy prediction sets in (15), up to the constant $1/(1-\gamma)$ that is irrelevant for the optimization. This shows that the non-fuzzy ‘as-if’ decision can indeed be viewed as a limiting case of the fuzzy procedure.
7 Covariates
In many applications, the observations $Z^{n+1}$ are decomposed into $Z^{n+1}=(X^{n+1},Y^{n+1})$, where the outcomes $Y^{n+1}$ come with covariates $X^{n+1}$. In such a setting, the covariate $X_{n+1}$ is observed alongside $Z^{n}=(X^{n},Y^{n})$, and we are to construct a prediction set for $Y_{n+1}$. In this section, we show how such covariates are easily incorporated into our framework. We only cover the non-fuzzy setting, as fuzzy prediction sets may be derived analogously.
When adding covariates, our prediction set $C_{\alpha}^{(Z^{n},X_{n+1})}$ now also depends on $X_{n+1}$. The coverage guarantee is also marginal over the covariates:
$P(Y_{n+1}\in C_{\alpha}^{(Z^{n},X_{n+1})})\geq 1-\alpha,$
for every $P\in\mathcal{P}$.
Following the discussion in Section 3.1, we may still construct a prediction set $C_{\alpha}$ for $Z^{n+1}$, and subsequently slice it at the realization $(Z^{n},X_{n+1})=(z^{n},x_{n+1})$ to obtain our prediction set $C_{\alpha}^{(z^{n},x_{n+1})}$ for $Y_{n+1}$:
$C_{\alpha}^{(z^{n},x_{n+1})}:=\{y\in\mathcal{Y}:C_{\alpha}(z^{n},(x_{n+1},y))=0\}.$
The optimality results remain the same, but we now maximize the power under the distribution
$Q=P_{*}^{Z^{n},X_{n+1}}\otimes\mu^{|(Z^{n},X_{n+1})}.$
7.1 Conformal prediction with covariates
In the application to the conformal setting, the only thing that changes is that we may use some additional structure in $Q$. In particular, we must replace $Q^{Z_{n+1}}=\mu^{|Z^{n}}$ by $Q^{Z_{n+1}}=P_{*}^{X_{n+1}|Z^{n}}\otimes\mu^{|(X_{n+1},Z^{n})}$
For example, the i.i.d. result in Proposition 2 now specializes to
$\textnormal{LR}^{|O(z^{n+1})}(z^{n+1})$ $\propto\frac{dQ^{Z_{n+1}}}{dQ^{Z_{1}}}(z^{n+1})$
$=\frac{d(P_{*}^{X_{n+1}}\otimes\mu^{|X_{n+1}})}{dP_{*}^{X_{n+1},Y_{n+1}}}(z_{n+1})$
$=\frac{dP_{*}^{X_{n+1}}\otimes d\mu^{|X_{n+1}}}{dP_{*}^{X_{n+1}}\otimes dP_{*}^{Y_{n+1}|X_{n+1}}}(z_{n+1})$
$=\frac{d\mu^{|X_{n+1}}}{dP_{*}^{Y_{n+1}|X_{n+1}}}(y_{n+1}),$
where the normalization constant is
$\frac{1}{(n+1)}\sum_{i=1}^{n+1}\frac{d\mu^{|X_{i}}}{dP_{*}^{Y_{n+1}|X_{i}}}(y_{i}).$
Here, we may write $P_{*}^{Y_{n+1}|X_{i}}$ instead of $P_{*}^{Y_{i}|X_{i}}$ due to the i.i.d. assumption.
To apply this, we may plug-in some estimator $\widehat{P}^{Y_{n+1}|X_{n+1}}$, for the kernel $P_{*}^{Y_{n+1}|X_{n+1}}$, or several estimators and average the resulting e-values. We use this in Section 8, where we estimate this kernel through a neural network.
8 Application: image recognition
In this section, we showcase our methodology in a classical Machine Learning problem: character recognition. Here, we use the same data and model as considered by gauthier2025values.
In particular, we consider the Federated Extended MNIST (FEMNIST) dataset (caldas2018leaf). In this dataset, $Y_{i}$ is one of 62 possible characters (a-z, A-Z, 0-9) and $X_{i}$ is a $28\times 28$ pixel image of a handwritten version of this character. We split the data into a training set (80%), calibration set (15%) and test set (5%), where the percentages are approximate because we put characters by the same writer into the same subset.
We use the training data to ‘train’ the same LeNet-inspired (lecun1998mnist) neural network as used by gauthier2025values, which is an estimator of the kernel $P_{*}^{Y_{n+1}|X_{n+1}}$. We plug this estimator into the likelihood ratio described in Section 7.1, and subsequently use this to construct several utility-optimal e-variables, as described in Section 5. The utilities we consider are all special cases of the capped power-utility $U:x\mapsto((x\wedge 1/\alpha)^{h}-1)/h$ framework introduced by koning2025continuoustestingunifyingtests, which interpolates between log-optimal e-value ($\alpha=0$ and $h\to 0$) and Neyman–Pearson-optimal e-value ($\alpha>0$ and $h=1$). We compute the fuzzy prediction set by taking the calibration set to be $Z^{n}$, and we select the image of a single observation from the test set to be $X_{n+1}$, so that our fuzzy prediction sets are on the true label $Y_{n+1}$.
In Figure 5, we plot the resulting fuzzy prediction sets for the label $Y_{n+1}$ of the image $X_{n+1}$ given in Figure 6. Here, each bar represents the amount of evidence (e-value) against every character. The plots are sorted from least-to-most evidence, on which they all agree (up to ties) by construction. Here, we see that the parameters $\alpha$ and $h$ are jointly able to change the shape of the fuzzy prediction set quite dramatically.
The approximate staircase-shape of the capped power-utility plot (bottom-left) makes it a good starting point to explain how we may interpret these plots. Ignoring the slightly elevated evidence at characters $G$, $b$ and $a$, this fuzzy prediction set can be interpreted as simultaneously reporting multiple prediction sets for different data-dependent level certificates $\widetilde{\alpha}$:
[…]
We may contrast this to the Neyman–Pearson plot (bottom-right), for which this becomes:
$C_{\widetilde{\alpha}}=\begin{cases}\{d,0,G,b\},&\textnormal{ for }0.001\leq\widetilde{\alpha},\\ \textnormal{all characters},&\textnormal{ for }\widetilde{\alpha}<0.001.\end{cases}$
This shows that a non-fuzzy prediction set yields a much more refined and non-binary expression of the available evidence. Comparing the two, a benefit of this fuzzy prediction set is that it reports a prediction set $\{d\}$ at a certificate of level $0.01$, which is unavailable for the Neyman–Pearson prediction set. The cost is that $\{d,0,G,b\}$ only comes with a certificate of $0.00125$, instead of the $0.001$ offered by the Neyman–Pearson variant.
The capped log-utility (top-right) and especially the uncapped log-utility (top-left) show a much wider spectrum of evidence. The uncapped log-utility displays strong evidence against characters such as $9$ and $R$, which is not visible in the other fuzzy prediction sets that are capped at $1000$. This may be of interest if the subsequent task is to eliminate certain characters from contention, instead of determining the correct character. Indeed, the log-utility maximizer reports at least 13 characters for post-hoc significance levels $\widetilde{\alpha}<0.1$, which does not seem helpful if we desire to identify the ‘correct’ character. This suggests that log-utility, which is often pushed as ‘the right utility function’ in the e-value literature, is not necessarily ideal for constructing fuzzy prediction sets. At the same time, such behavior may be attractive in different applications, such as a preliminary medical diagnosis to help rule out several implausible diseases.
An important open question is what kind of utility functions give rise to desirable fuzzy prediction sets for certain subsequent decision tasks. For now, the parameters $\alpha$ and $h$ seem like two useful instruments to shape our utility function, and thereby the resulting prediction sets: $\alpha$ caps the maximum evidence that we are interested in, and $h\in[-\infty,1]$ influences how ‘risky’ or ‘aggressive’ the e-variable is.
Figure 5: Expected-power-utility-optimal $0,1/\alpha]-$valued fuzzy prediction sets over possible labels, for $h=0$ (log-utility), $h=1/2$ (power-utility) and $h=1$ (Neyman–Pearson), for values $\alpha=0$ (top-left) and $\alpha=0.001$ (others). The true character label is marked in red (d).
Figure 6: Handwritten character used in the application.
9 Data availability
Code to replicate the application may be found at https://github.com/nickwkoning/fuzzy_prediction_sets.
10 Acknowledgements
We thank Etienne Gauthier, Peter Grünwald, Guneet Dhillon, Yash Nair and Junu Lee for their feedback and discussions.